OpenStudy (mary.rojas):

I got stuck with this problem,can someone help?? I will post the question and then my work so far

OpenStudy (mary.rojas):

Number 63

OpenStudy (solomonzelman):

which ones are you clear with and which ones are you troubled?

OpenStudy (mary.rojas):

I am not sure if that is the final answer

OpenStudy (mary.rojas):

Or is there another rule that could be applied here to shorten it

OpenStudy (solomonzelman):

(btw, I can't see the second column fully)

OpenStudy (mary.rojas):

Oh it is only the #63 problem Not the others

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle {\rm Ln } \left(\frac{x}{x-1}\right) +{\rm Ln } \left(\frac{x+1}{x}\right) -{\rm Ln } \left(x^2-1 \right) }\)

OpenStudy (mary.rojas):

I was just given a hint that helped me cancel out @SolomonZelman :) thank you

OpenStudy (mary.rojas):

I will post in a second!

OpenStudy (mary.rojas):

Okay,you will end up getting -2In(x-1) :) Thanks for the help

OpenStudy (solomonzelman):

yes, correct@!

OpenStudy (solomonzelman):

i didn't even help you with this much to be honest... you figured it out on your own... god job!

OpenStudy (mary.rojas):

Thank you though :) Actually mathmate gave me a hint to where I was stuck at :)